OpenStudy (anonymous):

a cube has a diagonal of 6. how long is one edge of a cube?

OpenStudy (anonymous):

36=2c^2 c = sqrt18 c =3sqrt2

OpenStudy (anonymous):

for some reason, the answer sheet says it's 2 square root of 3. :| is it wrong?

OpenStudy (anonymous):

wait a minute i did a lil wrong it's a cube i took it as square

OpenStudy (anonymous):

okay got it side = sqrt12 = 2sqrt3

OpenStudy (anonymous):

first the face diagonal = sqrt2xside then when we apply Pythagoras = 3c^2 =36 =>c=sqrt12

OpenStudy (anonymous):

i'm so sorry but i got lost in your explanation.

OpenStudy (anonymous):

okay actually this is 3d so it gets complicated but it's okay okay now face diagonal imagine a square the diagonal is sqrt2c if c is side of the cube

OpenStudy (anonymous):

now in a cube diagonal a facediagonal and a side are involved facediagonal is sqrt2c and the side is c that makes cube diagonal = sqrt(2c^2 +c^2) now cube diagonal is 6 hence 36 = 3c^2 c = sqrt12 c=2sqrt3

OpenStudy (anonymous):

now if your still unable to understand draw a diagram and then go through this whole process again